OpenStudy (dimples97):

Help me solve please! Determine the solution of the system. Show your work. Y = 5x – 9 Y = 2x + 6

pooja195 (pooja195):

Hi there! Welcome to OpenStudy :) Any ideas on where to start?

OpenStudy (asiannerd):

You can subtract the two equations by subtracting both sides to get the equations in the form ax+by=c

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @asiannerd You can subtract the two equations by subtracting both sides to get the equations in the form ax+by=c \(\color{#0cbb34}{\text{End of Quote}}\) There is an easier way

OpenStudy (asiannerd):

You can also you substitution by isolating one of the x's and plug in the x into the other equation

pooja195 (pooja195):

Since they already tell you what y equals Y = 5x – 9 <---equation 1 Y = 2x + 6 <----Equation 2 Take equation 1 and plug it into equation 2 Like this: 5x-9=2x+6 Now solve for x

OpenStudy (dimples97):

okay so you solve both equations for x then what?

pooja195 (pooja195):

It's one equation 5x-9=2x+6 You need to solve for x in this equation

OpenStudy (dimples97):

My bad got a little confused

OpenStudy (dimples97):

But thankyou both (:

OpenStudy (asiannerd):

So here if you subtract, you want (-5x+y=-9)-(-2x+y=6)

pooja195 (pooja195):

^not really D: 5x-9=2x+6  You need to come like terms

OpenStudy (dimples97):

so then it's 5x=2x15?

pooja195 (pooja195):

*5x=2x+15 Yep now continue

OpenStudy (dimples97):

hmmmm.. then divide 2 by 5??

pooja195 (pooja195):

No but we can subtract it :)

pooja195 (pooja195):

5-2=?

OpenStudy (dimples97):

3 ahah

pooja195 (pooja195):

Good 3x=15 any ideas on what's next?

pooja195 (pooja195):

^Remember that we are trying to get x ALONE.

OpenStudy (dimples97):

Ohhh okay, so now you still subtract right? 3 - 15?

pooja195 (pooja195):

hmm no quite...we would divide to get x alone 15/3=?

OpenStudy (dimples97):

or divide

pooja195 (pooja195):

*not

pooja195 (pooja195):

yep lol

OpenStudy (dimples97):

sooo x=5?

pooja195 (pooja195):

Yep yep! :) Now pick an equation, equation 1 or 2 ? (just like pick a card any card :P )

OpenStudy (dimples97):

hmmm equation 1? :o

pooja195 (pooja195):

okie good =) We know that x=5 so all we do is plug 5 into equation 1 to get our y value Y = 5(5) – 9 Solve for y

OpenStudy (dimples97):

y=25-9?

pooja195 (pooja195):

yep keep going

OpenStudy (dimples97):

y= hmmm 25?

pooja195 (pooja195):

not quite! 25-9=?

OpenStudy (dimples97):

Ohhh! sorry i completely forgot the 9

OpenStudy (dimples97):

y=16!

pooja195 (pooja195):

No worries haha yep thats right So we have (5,16) where x=5 and y=16

OpenStudy (dimples97):

And thats it?

pooja195 (pooja195):

Did I make it too easy? ;)

OpenStudy (dimples97):

Hahah you helped me alot (: sorry for being a little difficult!

pooja195 (pooja195):

Lol you're fine no worries! And you're welcome ^_^